Minute of KAGRA commissioning meeting
- Date: 2018/8/7 11AM~12PM
- Attendance: Miyakawa, Aso, Yokozawa
- eZuce: Miyoki, Suzuki, Kimura, Enomoto, Ushiba, Yamamoto, Inoue, Uchiyama
Check list for IXC
- Meeting will be assigned on one of 8/16 12:30
- Frost of nitrogen at view port inside of chamber of ETMX was seen by leakage.
- ETMY frost was another issue.
- Number of stored channels should be checked.
- Some BIOs are not working. DC-DC power supply should be tested to fix them.
- PSL sensors have been installed.
- Please send requests for daily page to PEM group.
- Camera of Tcam will be replaced.
- They will be used from the middle of September.
- Next meeting: 8/21.